# http://www.sharif.edu

### v1

@@ Abstract
CLIP models have recently shown to exhibit Out of Distribution (OoD) generalization capabilities. However, Compositional Out of Distribution (C-OoD) generalization, which is a crucial aspect of a model’s ability to understand unseen compositions of known concepts, is relatively unexplored for the CLIP models. Our goal is to address this problem and identify the factors that contribute to the C-OoD in CLIPs. We noted that previous studies regarding compositional understanding of CLIPs frequently fail to ensure that test samples are genuinely novel relative to the CLIP training data. To this end, we carefully synthesized a large and diverse dataset in the single object setting, comprising attributes for objects that are highly unlikely to be encountered in the combined training datasets of various CLIP models. This dataset enables an authentic evaluation of C-OoD generalization. Our observations reveal varying levels of C-OoD generalization across different CLIP models. We propose that the disentanglement of CLIP representations serves as a critical indicator in this context. By utilizing our synthesized datasets and other existing datasets, we assess various disentanglement metrics of text and image representations. Our study reveals that the disentanglement of image and text representations, particularly with respect to their compositional elements, plays a crucial role in improving the generalization of CLIP models in out-of-distribution settings. This finding suggests promising opportunities for advancing out-of-distribution generalization in CLIPs. For more details and access to our dataset, please visit our project GitHub repository.
@@ 2.1.2 Selection of Attributes (Adjectives)
We then selected 140 adjectives from the Visual Attributes Words (VAW) dataset[16]. These adjectives span various categories, including color, material, and texture, allowing us to create a wide range of descriptive combinations for image generation. A complete list of the 140 adjectives used from the Visual Attributes Words (VAW) dataset is provided in 9Appendix 7.2.
@@ 2.1.4 Filtering Process
Step 3 - Verification of OoD Status: The final step in our filtering process was to ensure the OoD nature of our dataset. We used the Faiss library for a K-nearest neighbors search to compare our generated images against those in the LAION, CommonPool, YFCC, and CC datasets. Images were considered unique and retained in our dataset if no closely matching analogs were found, based on human evaluation. This rigorous approach ensured the novelty and uniqueness of our dataset by excluding combinations that had similar matches in the referenced datasets.
@@ 3 Comparison of CLIP Models on ImageNet-AO
To evaluate the CLIP model performance in the classification tasks, we adopted the evaluation method developed by [17], similar to the zero-shot evaluation approach described in [2]. Our evaluation involves providing the model with the actual images and various captions, obtaining embeddings for both the images and texts, and calculating their cosine similarities. This allows us to estimate the relevance of the captions to the image content, similar to a classification task. Since we only had class labels (attribute-object pairs) for the images, and many captions could match the class label, 80 captions were created for each class using several templates, similar to the approach described in [2]. We prepared embeddings for these captions and took their average to obtain a final embedding, which was used in the zero-shot evaluation. For the test sets, all 1000 classes of ImageNet were used as the in-distribution set and expanded the number of classes to approximately 23000 for the OoD set. The CLIP evaluations are shown in Fig. 1.
@@ 5.2 Disentanglement of Fine-Grained Factors
In the field of Disentanglement Representation Learning, the concept of disentanglement is explored from two distinct perspectives: fine-grained factors at the dimension level and coarse-grained factors at the vector level [26]. Our initial investigation into CLIP models, utilizing our curated dataset, provided insights into coarse-grained disentanglement (e.g., separating attributes and objects as two factors) and revealed multifaceted evaluation metrics. Moving forward, we aim to delve into the realm of fine-grained disentanglement at the dimension level. However, our current dataset poses inherent limitations in segregating factors at such a granular level. Consequently, to facilitate a comprehensive evaluation of fine-grained disentanglement, it becomes necessary to adopt specialized datasets designed explicitly for disentanglement studies within this domain.
Figure 7: Disentanglment datasets. a: Sprites dataset, consist of 6 factor and 54,000 images b: Shape3D, consist of 5 factor and 32,000 images
@@ 5.2.1 More Analysis on decomposability of the representation space
LAION ViT-L/14 63.47% 2 40
LAION ViT-B/16 60.01% 5 60
LAION ViT-B/32 59.41% 7 90
OpenAI ViT-L/14 50.32% 3 5
OpenAI ViT-B/16 47.50% 4 10
OpenAI ViT-B/32 45.8% 6 30
CC RN50 25.70% 15 200
YFCC RN50 11.50% 21 250

### v2

@@ Abstract
CLIP models have recently shown to exhibit Out of Distribution (OoD) generalization capabilities. However, Compositional Out of Distribution (C-OoD) generalization, which is a crucial aspect of a model’s ability to understand unseen compositions of known concepts, is relatively unexplored for the CLIP models. Our goal is to address this problem and identify the factors that contribute to the C-OoD in CLIPs. We noted that previous studies regarding compositional understanding of CLIPs frequently fail to ensure that test samples are genuinely novel relative to the CLIP training data. To this end, we carefully synthesized a large and diverse dataset in the single object setting, comprising attributes for objects that are highly unlikely to be encountered in the combined training datasets of various CLIP models. This dataset enables an authentic evaluation of C-OoD generalization. Our observations reveal varying levels of C-OoD generalization across different CLIP models. We propose that the disentanglement of CLIP representations serves as a critical indicator in this context. By utilizing our synthesized datasets and other existing datasets, we assess various disentanglement metrics of text and image representations. Our study reveals that the disentanglement of image and text representations, particularly with respect to their compositional elements, plays a crucial role in improving the generalization of CLIP models in out-of-distribution settings. This finding suggests promising opportunities for advancing out-of-distribution generalization in CLIPs. For more details and access to our dataset, please visit https://github.com/abbasiReza/CLIP-COoD.
@@ 2.1.2 Selection of Attributes (Adjectives)
We then selected 140 adjectives from the Visual Attributes Words (VAW) dataset[16]. These adjectives span various categories, including color, material, and texture, allowing us to create a wide range of descriptive combinations for image generation. A complete list of the 140 adjectives used from the Visual Attributes Words (VAW) dataset is provided in Appendix 7.3.
@@ 2.1.4 Filtering Process
Step 3 - Verification of OoD Status: The final step in our filtering process was to ensure the OoD nature of our dataset. We used the Faiss library [17] for a K-nearest neighbors search to compare our generated images against those in the LAION, CommonPool, YFCC, and CC datasets. Images were considered unique and retained in our dataset if no closely matching analogs were found, based on human evaluation. This rigorous approach ensured the novelty and uniqueness of our dataset by excluding combinations that had similar matches in the referenced datasets.
@@ 3 Comparison of CLIP Models on ImageNet-AO
To evaluate the CLIP model performance in the classification tasks, we adopted the evaluation method developed by [18], similar to the zero-shot evaluation approach described in [2]. Our evaluation involves providing the model with the actual images and various captions, obtaining embeddings for both the images and texts, and calculating their cosine similarities. This allows us to estimate the relevance of the captions to the image content, similar to a classification task. Given that our dataset only provided class labels (attribute-object pairs) for images, we expanded on this by creating 80 captions per class using various templates. This approach, inspired by the methodology described in [2], allows for a more comprehensive representation of each class. We generated embeddings for these captions and averaged them to produce a final embedding for each class, which was then used in our zero-shot evaluation. For the test sets, all 1000 classes of ImageNet were used as the in-distribution set and expanded the number of classes to approximately 21000 for the OoD set. The CLIP evaluations are shown in Fig. 1.
@@ 5.2 Disentanglement of Fine-Grained Factors
In the field of Disentanglement Representation Learning, the concept of disentanglement is explored from two distinct perspectives: fine-grained factors at the dimension level and coarse-grained factors at the vector level [27]. Our initial investigation into CLIP models, utilizing our curated dataset, provided insights into coarse-grained disentanglement (e.g. separating attributes and objects as two factors) and revealed multifaceted evaluation metrics. Moving forward, we aim to delve into the realm of fine-grained disentanglement at the dimension level. However, our current dataset poses inherent limitations in segregating factors at such a granular level. Consequently, to facilitate a comprehensive evaluation of fine-grained disentanglement, it becomes necessary to adopt specialized datasets designed explicitly for disentanglement studies within this domain.
Figure 7: Disentanglement datasets. a: Sprites dataset, consist of 6 factor and 54,000 images b: Shape3D, consist of 5 factor and 32,000 images
@@ 5.2.1 More Analysis on decomposability of the representation space
LAION ViT-L/14 64.61% 2 40
LAION ViT-B/16 61.55% 5 60
LAION ViT-B/32 61.05% 7 90
OpenAI ViT-L/14 52.28% 3 5
OpenAI ViT-B/16 49.22% 4 10
OpenAI ViT-B/32 47.07% 6 30
CC RN50 26.64% 15 200
YFCC RN50 12.23% 21 250